A balloon filled with hydrogen gas had a volume of 45000 L at a pressure of 760 mmHg and a temperature of 0 ∘C. When the balloon reached an altitude of 1000 m, the pressure was 608 mmHg and the temperature was − 9 ∘C. What was the volume in liters of the balloon at these conditions, if the amount of gas does not change?
Answer #1

From gas laws, P1V1/T1 = P2V2/T

Initial conditions:   P1 = 760 mmHg = 1 atm, V1 = 45000 L, T1 = 0 ∘C = 273 K

Final conditions:   P2 = 608 mmHg = 0.8 atm, V1 = ? L, T1 = -9 ∘C = 264 K

Therefore, 1*45000/273 = 0.8*V2/264 = 54395.60 L

The final volume in liters of the balloon = 54395.60 L